The base of a triangle is 6cm greater than twice the height. The area is 28cm squared. What is the height and base?
1
Expert's answer
2020-04-11T17:39:43-0400

"S=\frac{1}{2}(2\times h+6)\times h=28;
h^2 +3\times h=28; h^2 +3\times h-28=0;D=9+4\times 28 = 121;
h = \frac{-3+11}{2};h=4;2\times h+6=8+6=14;S=\frac{1}{2}(14\times 4)=28;"

Answer: the height is 4cm and base is 14cm